OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = -x - 3? A) y = -x - 1 B) y = -x + 5 C) y = x + 1 D) y = x + 5

OpenStudy (hoblos):

-1*slope =-1 => slope = 1 => y-3=x-2 => y=x+1

OpenStudy (anonymous):

thnks